(a) State and explain the two traditional postulates from which special relativity originates. (b) Draw a plot of the Lorentz factor for 0 ≤ β ≤ 0.9 to see how its value changes.


1
Expert's answer
2021-06-07T09:32:30-0400

(a)    1. The laws of physics are invariant (that is, identical) in all inertial frames of reference (that is, frames of reference with no acceleration). 2. The speed of light in vacuum is the same for all observers, regardless of the motion of the light source or observer.


(b) Lorentz factor is "\displaystyle \gamma = \frac{1}{\sqrt{1-\beta^2}}" where "\beta = v/c" .